OpenStudy (anonymous):

If x = -8 find the matching value for y. -2x-4y=8

OpenStudy (anonymous):

I need the ANSWER please.

OpenStudy (cwrw238):

-2(-8) - 4y = 8 -4y = 8 -16 -4y = -8 4y = 8 now divide by 4 and you have it

OpenStudy (anonymous):

Thank you :)

OpenStudy (cwrw238):

yw